OpenStudy (anonymous):

*.*.*.*.*.*I will give medal*.*.*.*.*.* How can sine's rule be used in this???

OpenStudy (e.mccormick):

You mean: For any triangle ABC, \(\dfrac{\sin A}{a}=\dfrac{\sin B}{b}=\dfrac{\sin C}{c}\) ?

OpenStudy (anonymous):

 Find all letters.

OpenStudy (e.mccormick):

Start with the Law of Cosines.

OpenStudy (anonymous):

no i need sine rule

OpenStudy (e.mccormick):

You have two unknowns that way. You need to start with Cosines to get it to one unnown, then use sines.

OpenStudy (anonymous):

ok so can u help me out

OpenStudy (e.mccormick):

So, do you know the Law of Cosines?

OpenStudy (anonymous):

\[c ^{2}=a ^{2}+b ^{2}-2ab \cos C\] \[c=\sqrt{(7^{2}+8^{2}-2(7)(8)\cos 50)}\] \[c=6.404\] is this right

OpenStudy (e.mccormick):

I got 6.40373232694, so yah.

OpenStudy (anonymous):



OpenStudy (anonymous):

ok now i can do it thanks.

OpenStudy (e.mccormick):

Now you can use that and sines to find \(\angle A\).